OpenStudy (anonymous):

Find the distance between -2/5 and -7

OpenStudy (gw2011):

-7=-35/5 The difference between -2/5 and -35/5=-33/5=-6.6 OR -2/5=-0.4 Again, find the difference between -0.4 and -7=-6.6